# Synergistic Multiscale Detail Refinement via Intrinsic Supervision for Underwater Image Enhancement

### 15 Dec 2023 v3

@@ Introduction
Enhancing low-quality images poses significant challenges to the field of computer vision (Ma et al. 2023; Khan et al. 2023). These challenges arise primarily from scattering and blurring effects unique to aquatic environments, which inherently manifest in a multi-scale manner (Zhou et al. 2023b). Particulate matter and water turbulence at different scales have different effects on different scales have different effects on different parts of an image, leading to the degradation of multi-scale correlated features. The underwater image formation model (UIFM) is represented as:
@@
The core objective of underwater image enhancement (UIE) is to accurately estimate $I$ from Eq. (1) (Zhou et al. 2023a). This is crucial because degradation levels in underwater images can differ across pixels due to variations in distance. There are many existing enhancement techniques (Zhuang, Li, and Wu 2021), which either globally or locally extract features (Liu et al. 2022a) (Liu et al. 2023). These methods often leverage filters (Zhou et al. 2023c) and color correction techniques (Zhou et al. 2022) to improve image quality. However, as illustrated in Figure 1, the inherent scale-related characteristics in underwater scenes (Jiang et al. 2020)(Hou et al. 2023) demand attention. Pixels at similar depths can exhibit consistent degradation levels. By recognizing these depth-informed degradation patterns, we can gain a richer understanding of the structure of the scene. In light of this, we proposed Synergistic Multiscale Detail Refinement via Intrinsic Supervision for Underwater Image Enhancement (SMDR-IS). This approach uniquely leverages low-resolution images as auxiliary inputs, providing additional insights into scene degradation.
SMDR-IS simultaneously harnesses multiple degradations to mitigate the loss of detailed features resulting from multiple feature extractions. The low-degradation stages serve to counteract the excessive abstraction from redundant extractions, ensuring that the intrinsic relationship between features and the original image remains intact.
While feature extraction at the original yields high-level features, which can often capture broad data characteristics at the expense of task-specific critical details. By integrating low-degradation images and features from the initial layers, both credibility and applicability are enhanced. Furthermore, the incorporation of ASISF guarantees that irrelevant information is excluded from the original resolution.
@@
(1) Addressing the limited of scale-related features in current underwater image enhancement methods, which often results in incomplete restoration of scene details, we propose a synergistic multiscale detail refinement via intrinsic supervision for underwater image enhancement.
(2) We design a new attention, namely Bifocal Intrinsic-Context Attention, based on a dual-path approach, ensuring that both feature enhancement and contextual semantic information are addressed. Additionally, we seamlessly integrate resolution-guided supervision to boost computational efficiency without sacrificing detail enhancement quality.
(3) With the aim of refining low-resolution feature information, we introduce the Adaptive Selective Intrinsic Supervised Feature (ASISF) module. This regulates feature propagation, enhancing image quality and avoiding blurring caused by utilizing the overlay of multiscale scene details.
@@
Prior-based methods rely on explicit degradation models or pre-existing knowledge to calibrate model parameters. For instance, UColor (Li et al. 2021) which is guided by transmission, employs a transmission-driven image enhancement network using GDCP (Peng, Cao, and Cosman 2018). (Zhou et al. 2023d) uses an augmented U-Net to fuse inputs, such as the original, color-corrected, and contrast-enhanced images, to effectively leverage features. Unsupervised underwater image restoration method (USUIR) (Fu et al. 2022) involves designing a transmission subnet, a global background subnet, and a scene radiance subnet to estimate parameters of UIFM, facilitating Photo-realistic image restoration. Zhang et al. (Zhang et al. 2023) proposed Rex-Net, which applied the Retinex theory to enhance underwater images. (Mu et al. 2023) employs physical knowledge to design the Adaptive Transmission-Guided Module to guide the network. While these techniques have shown promising results, the intricate and unpredictable nature of underwater environments sometimes undermines the efficacy of the priors used, potentially limiting model adaptability.
@@
Our proposed model integrates four stages of multi-resolution image inputs, equipping the network and the related features with diverse scale scene information of the input image. Initially, we employ downsampling on the original image to derive underwater scenes of different scales. While prevailing UNet-based underwater enhancement methods (Li et al. 2021) typically utilize three downsampling steps, our approach extends this to four stages. This can capture a richer array of scale scene information, comprising the original image and its three downscaled images. To harness the scale-correlated attributes of image scenes, SMDR-IS features a resolution encoder, further supplemented by three stages for lower-resolution images.
Each encoder stage incorporates the Bifocal Intrinsic-Context Attention Module (BICA) to amplify feature details. BICA can adeptly fuse global and local features, optimizing image enhancement. With the aid of resolution-guided supervision, BICA strikes a harmonious balance between computational efficiency and details enhancement.
To bolster the utility of low-resolution features, we proposed the Adaptive Selective Intrinsic Supervised Feature (ASISF) module. This module governs feature propagation, elevating image enhancement outcomes, and curbing information blurring that may arise from superimposing multi-scale scene details.
@@
In the second branch, acknowledging the pivotal role of spatial contextual relationships in underwater image restoration, we design the Hierarchical Context-Aware Feature Extraction (HCAFE) module (detailed in the supplementary material). This module functions within the original feature domain, extracting image representation features through diverse receptive fields.
BCIA employs innovative dual-path attention, integrating both channel and spatial attention to facilitate enhanced extraction and fusion of both global and local image features. Leveraging a resolution-guided supervision approach, BCIA achieves superior computational efficiency without sacrificing detail quality. Its unique fusion of global context with local nuances, along with its adaptability to existing models, positions BCIA as a groundbreaking advancement in UIE.
Resolution-Guided Intrinsic Attention Module Underwater image optimization is heavily influenced by the contextual details of adjacent pixels. However, to manage network complexity, enhancement methods often utilize compact $3\times 3$ convolutional kernels for feature extraction. Although $3\times 3$ kernels are computationally efficient, the small receptive field hinders the network to capture extensive contextual nuances.
To address the limitation and broaden the receptive field, we propose the Resolution-Guided Intrinsic Attention Module (ReGIA), which enhances the preliminary features extracted from the input image by CFA. ReGIA is tailored to discern feature weight data in a lower-resolution latent space with a wilder receptive field. By expanding the receptive field of convolutional kernels, ReGIA is a guiding beacon that not only enhances the correlation of features in the original domain, but also learns features from its larger contextual receptive field.
@@
ASISF is meticulously designed to retain only the most pertinent features for processing in the original-resolution stage. When learning features from the low-resolution encoder, the features from the original resolution act as a reference. Specifically, within the encoder, the output of SMDR-IS is employed as a reference to guide the operation of the low-resolution encoder-decoder. Additionally, SMDR-IS harnesses multi-degradation GT, comparing the output of SMDR-IS with multi-degradation images, to instill intrinsic supervision. This strategy ensures that both the encoder and decoder components of multi-degradation align with the genuine degradation traits inherent in underwater images. Such alignment is pivotal for effective learning and generalization of the network, ultimately enhancing image restoration capabilities.
This is particularly significant in the context of image restoration, where the presence of noise, distortions, and varying levels of degradation can complicate the restoration process. By employing intrinsic supervision, ASISF effectively identifies and retains features meaningful for the restoration task, thereby improving the precision and efficiency of the restoration process.
@@
Unlike the traditional UNet, the multi-degradation decoder in SMDR-IS is uniquely designed to complement its encoder counterpart. Low-degradation information captures the degraded nuances of input images, which often stem from factors, such as noise, blurring, or compression artifacts. By integrating this information, the decoder can gain a deeper insight into these degradation patterns.
The fusion of low-degradation and residual information from the encoder equips the decoder with a holistic grasp of the complicated image degradation. This comprehensive understanding empowers the network to produce restoration results that are both precise and context-rich. Essentially, the low-degradation information acts as a guiding beacon, enhancing the decoder’s restoration capabilities. Consequently, the output images not only exhibit crisp details but also adeptly rectify various types of degradation. It is worth mentioning that within the $FR_{iD}$, the ASISF performs intrinsic-supervised feature selection on the output features when feeding data to the original degradation encoder. This ensures that only the most relevant feature is propagated through the image processing pipeline. In essence, SMDR-IS’s decoder uses low-degradation cues to adeptly counteract and rectify diverse image degradations, thereby notably improving the fidelity of the restored images.
@@
To enhance the precision of the restoration process, SMDR-IS employs a multi-degradation loss function to govern different modules of the network.
@@
Our method was implemented using PyTorch, with an NVIDIA Tesla V100 GPU, Intel(R) Xeon(R) Silver 4114 CPU, and 32GB RAM. For training, images were randomly cropped to a resolution of 256$\times$256, and we used a batch size of 44 and a learning rate of 0.0002. To ensure that SMDR-IS can generate outputs consistent with the original image size during testing, techniques, such as border padding were employed.
@@
To avoid any influence of color on the metrics (e.g., the impact observed with ULAP in Figure 4), we excluded the color component from CCF. As shown in Figure 4, SMDR-IS performs better than other methods in terms of visual results and performance metrics. Although ULAP, IBLA, and GDCP achieve commendable restoration results, their generalization capabilities are hindered by their dependence on priors. WaterNet and UColor display enhanced robustness but remain susceptible to extraneous inputs. Although FUnIE-GAN and UWCNN can achieve real-time efficiency, their expressive capability is somewhat constrained due to limited parameters. Although UDA and U-shape can both achieve the goal of underwater image enhancement, SMDR-IS combines scenes from multiple scales, providing better robustness for diverse underwater scenarios.
Table 1 tabulates the performance of different methods on the tree datasets, in terms of different performance metrics. In addition to the individual performance metrics, we also combine them to form a comprehensive metric, denoted as ”ALL”, which is the net sum of the $\uparrow$ values minus the $\downarrow$ values. This measurement can provide a holistic assessment by weighing multiple criteria. Remarkably, SMDR-IS achieves the best performance, in terms of the ALL score. Collectively, these results show the ability of SMDR-IS to adeptly restore intricate scenes, underpinned by the synergy from multiscale detail extraction and intrinsic supervision.
@@
Table 2 presents the comprehensive analysis of efficiency, performance (the Quality is the ”ALL” on UIEB Val in Table 1), and Aggregative (Agg) score. Although SMDR-IS does not obtain the highest efficiency, it still achieves a commendable frame rate of 16.4744 FPS, satisfying real-time demands. Moreover, we introduce the Agg score to provide a holistic performance evaluation, which considers both efficiency and performance, specifically, Agg = Quality-Time. As observed from Table 2, SMDR-IS achieves the highest score in Agg, which demonstrates the suitability of SMDR-IS for advanced underwater vision tasks.
@@ Conclusion
In this study, we propose a novel method for underwater image restoration, namely SMDR-IS, which can proficiently capture multi-scale scene information, by multi-resolution detail extraction with intrinsic supervision, and utilizing the inherent scale-related features in image scenes. To ensure optimal assimilation of information across different scales, we integrated low-resolution inputs within the original-resolution input, thereby amplifying the conveyance of scene information. To mitigate unnecessary interference from irrelevant scenes, we introduced ASISF, which is meticulously designed to regulate the feature propagation process. Furthermore, our multi-degradation loss function strategically guides the network during training. Optimization and constraints at each stage enhance the network’s ability to leverage information at different scales. In essence, SMDR-IS adeptly exploits multi-scale nuances to enrich scene information in underwater imagery. In the future, we will explore the integration of SMSR-IS into broader computer vision applications, like underwater robotics and autonomous underwater vehicles.
@@ Acknowledgments
Natural Science Foundation of China (Nos. 62301105 and 61702074), the Liaoning Provincial Natural Science Foundation of China (No. 20170520196), and the Fundamental Research Funds for the Central Universities (Nos. 3132019205 and 3132019354), National Key Research and Development Program of China (No.2018AAA0100400), China Postdoctoral Science Foundation (NO.2021M701780), and the Cultivation Program for the Excellent Doctoral Dissertation of Dalian Maritime University. We are also sponsored by CAAI-Huawei MindSpore Open Fund and the High Performance Computing Center of Dalian Maritime University.

### 18 Jan 2024 v4

@@ Introduction
Enhancing low-quality images poses significant challenges to the field of computer vision (Ma et al. 2023; Khan et al. 2023; Zhou et al. 2024). These challenges arise primarily from scattering and blurring effects unique to aquatic environments, which inherently manifest in a multi-scale manner (Zhou et al. 2023b). Particulate matter and water turbulence at different scales have different effects on different scales have different effects on different parts of an image, leading to the degradation of multi-scale correlated features. The underwater image formation model (UIFM) (Zhou et al. 2023e) is represented as:
@@
The core objective of underwater image enhancement (UIE) is to accurately estimate $I$ from Eq. (1) (Zhou et al. 2023a). This is crucial because degradation levels in underwater images can differ across pixels due to variations in distance. There are many existing enhancement techniques (Zhuang, Li, and Wu 2021), which either globally or locally extract features (Liu et al. 2022a) (Liu et al. 2023). These methods often leverage filters (Zhou et al. 2023c) and color correction techniques (Zhou et al. 2022) to improve image quality. As illustrated in Figure 1, similar degradation scenarios exist in different resolution stages, Therefore, the inherent scale-related characteristics in multi-scale underwater scenes (Jiang et al. 2020)(Hou et al. 2023) demand attention. By recognizing inherent scale-related degradation patterns, we can gain a richer understanding of the structure of the scene. In light of this, we proposed Synergistic Multiscale Detail Refinement via Intrinsic Supervision for Underwater Image Enhancement (SMDR-IS). SMDR-IS uniquely leverages low-resolution images as auxiliary inputs, providing additional insights into scene degradation.
SMDR-IS simultaneously harnesses multiple degradations to mitigate the loss of multiple scene features resulting from feature extractions. While feature extraction at the original yields high-level features produce high-level features, it is often possible to capture finer data features at the expense of larger-scale scene information. The low-degradation stages serve to counteract the excessive abstraction from redundant extractions, ensuring that the intrinsic relationship between features and the original image remains intact. By integrating low-degradation images and features from the initial layers, both credibility and applicability are enhanced. Furthermore, the incorporation of ASISF guarantees that irrelevant information is excluded from the original resolution.
@@
(1) Addressing the limitations of scale-related features in current underwater image enhancement methods, which lead to incomplete restoration of scene details, we propose a synergistic multiscale detail refinement via intrinsic supervision for underwater image enhancement.
(2) We design a new attention, namely Bifocal Intrinsic-Context Attention, based on a dual-path approach, ensuring that both feature enhancement and contextual semantic information are addressed. Additionally, we integrate resolution-guided supervision to boost computational efficiency without sacrificing detail enhancement quality.
(3) With the aim of refining low-resolution feature information, we introduce the Adaptive Selective Intrinsic Supervised Feature (ASISF) module. ASISF regulates feature propagation, enhancing image quality and avoiding blurring caused by utilizing the overlay of multiscale scene details.
@@
Prior-based methods rely on explicit degradation models or pre-existing knowledge to calibrate model parameters. For instance, UColor (Li et al. 2021) which is guided by transmission, employs a transmission-driven image enhancement network using GDCP (Peng, Cao, and Cosman 2018). (Zhou et al. 2023d) uses an augmented U-Net to fuse inputs, such as the original, color-corrected, and contrast-enhanced images, to effectively leverage features. Unsupervised underwater image restoration method (USUIR) (Fu et al. 2022) involves designing a transmission subnet, a global background subnet, and a scene radiance subnet to estimate parameters of UIFM, facilitating Photo-realistic image restoration. Zhang et al. (Zhang et al. 2023) proposed Rex-Net, which applied the Retinex theory to enhance underwater images. (Mu et al. 2023) employs physical knowledge to design the Adaptive Transmission-Guided Module to guide the network. While these methods have shown promising results, the intricate and unpredictable nature of underwater environments sometimes undermines the efficacy of the proposed priors, potentially limiting model adaptability.
@@
Our proposed model integrates four stages of multi-resolution image inputs, equipping the network and the related features with diverse scale scene information of the input image. Initially, we employ downsampling on the original image to derive underwater scenes of different scales. Since prevailing UNet-based underwater enhancement methods (Li et al. 2021) typically utilize three downsampling steps, we adopt three downscaled stages corresponding to each stage of the UNet. This can capture a richer array of scale scene information. To harness the scale-correlated attributes of image scenes, SMDR-IS proposed the multi-degradation encoder, further supplemented by three stages for lower-resolution images.
Each encoder stage incorporates the Bifocal Intrinsic-Context Attention Module (BICA) to extract feature details. BICA can adeptly fuse global and local features, optimizing image enhancement. With the aid of resolution-guided supervision, BICA strikes a harmonious balance between computational efficiency and details enhancement. To bolster the utility of low-resolution features, we proposed the Adaptive Selective Intrinsic Supervised Feature (ASISF) module. This module governs feature propagation, elevating image enhancement outcomes, and curbing information blurring that may arise from superimposing multi-scale scene details.
@@
In the second branch, acknowledging the pivotal role of spatial contextual relationships in underwater image restoration, we design the Hierarchical Context-Aware Feature Extraction (HCAFE) module (detailed in the supplementary material). This module functions within the original feature domain, extracting image representation features through hierarchical context attention.
BCIA leverages a resolution-guided supervision approach, achieving superior computational efficiency without sacrificing detail quality. BCIA employs an innovative dual-path attention mechanism, comprehensively addressing both the influence of neighboring pixel regions and spatial contextual relationships.
Resolution-Guided Intrinsic Attention Module Underwater image optimization is heavily influenced by the contextual details of adjacent pixels. However, to reduce the computational efficiency of the network, enhancement methods often utilize compact $3\times 3$ convolutional kernels for feature extraction. Although $3\times 3$ kernels are computationally efficient, the small receptive field hinders the network to capture extensive contextual features.
To address the limitation and broaden the receptive field, we propose the Resolution-Guided Intrinsic Attention Module (ReGIA), which enhances the preliminary features extracted from the input image by CFA. ReGIA is tailored to discern feature weight data in a lower-resolution latent space with a large feature receptive field. Guided by the lower-resolution latent space, ReGIA serves as a guiding beacon that not only enhances the correlation of features in the original domain but also strives to improve the computational efficiency of the network.
@@
ASISF is meticulously designed to retain only the most pertinent features to supplement image feature extraction and reconstruction in the original resolution stage. When learning features from the low-resolution encoder, the features from the original resolution act as a reference. This strategy ensures that both the encoder and decoder components of multi-degradation align with the genuine degradation traits inherent in underwater images.
This is particularly significant in the context of image restoration, where the presence of noise, distortions, and varying levels of degradation can complicate the restoration process. By employing intrinsic supervision, ASISF effectively identifies and retains features meaningful for the restoration task, thereby improving the precision and efficiency of the restoration process. The feature adaptive selection is pivotal for effective learning and generalization of the SMDR-IS, ultimately enhancing image restoration capabilities by supplementing inherent scale-related degradation patterns at multi-resolution.
@@
Unlike the traditional UNet, the multi-degradation decoder in SMDR-IS is uniquely designed to complement its encoder counterpart. Low-degradation information captures the inherent scale-related characteristics of input images. By integrating this information, the decoder can gain a deeper insight into these degradation patterns and multiple scenes.
The fusion of low-degradation and residual information from the encoder equips the decoder with a holistic grasp of the complicated image degradation. This comprehensive understanding empowers the network to produce restoration results that are both precise and context-rich. Essentially, the low-degradation information acts as the supplementary branch, enhancing the decoder’s restoration capabilities. Consequently, the output images not only exhibit crisp details but also accurately recover information from multiple scenes. It is worth mentioning that within the $FR_{iD}$, the ASISF performs intrinsic-supervised feature selection on the output features when feeding data to the original degradation encoder. This ensures that only the most relevant feature is propagated through the image processing pipeline. In essence, the multi-degradation decoder of SMDR-IS uses low-degradation cues to ingeniously supplement inherent scale-related characteristics, thereby notably improving the fidelity of the restored images.
@@
To enhance the precision of the restoration process, SMDR-IS employs a multi-degradation loss function to govern different modules of the network, which compares the multi-resolution enhancement image of SMDR-IS with multi-resolution ground truth, to instill multi-resolution supervision.
@@
Our method was implemented using PyTorch, with an NVIDIA Tesla V100 GPU, Intel(R) Xeon(R) Silver 4114 CPU, and 32GB RAM. For training, images were randomly cropped to a resolution of 256$\times$256, and we used a batch size of 44 and a learning rate of 0.0002. To ensure that SMDR-IS can generate outputs consistent with the original image size during testing, we adopted the border padding techniques.
@@
To avoid any influence of color on the metrics (e.g., the impact observed with ULAP in Figure 4), we excluded the color component from CCF. As shown in Figure 4, SMDR-IS performs better than other methods in terms of visual results and performance metrics. Although ULAP, IBLA, and GDCP achieve commendable restoration results, their generalization capabilities are hindered by their dependence on priors. WaterNet and UColor display enhanced robustness but remain susceptible to extraneous inputs. Although FUnIE-GAN and UWCNN can achieve real-time efficiency, their expressive capability is somewhat constrained due to limited parameters. Although UDA and U-shape can both achieve the goal of underwater image enhancement, SMDR-IS combines the inherent scale-related features from multiple scales, providing better robustness for diverse underwater scenarios.
Table 1 tabulates the performance of different methods on the three datasets, in terms of different performance metrics. In addition to the individual performance metrics, we also combine them to form a comprehensive metric, denoted as ”ALL”, which is the net sum of the $\uparrow$ values minus the $\downarrow$ values. This measurement can provide a holistic assessment by weighing multiple criteria. Remarkably, SMDR-IS achieves the best performance, in terms of the ALL score. Collectively, these results show the ability of SMDR-IS to adeptly restore intricate scenes, underpinned by the synergy from multiscale detail extraction and intrinsic supervision.
@@
Table 2 presents the comprehensive analysis of efficiency, performance (the Quality is the ”ALL” on UIEB Val in Table 1), and Aggregative (Agg) score. Although SMDR-IS does not obtain the highest efficiency, it still achieves a commendable frame rate of 16.4744, satisfying real-time demands. Moreover, we introduce the ’Agg’ score to provide a holistic performance evaluation, which considers both efficiency and performance, specifically, Agg = Quality-Time. As observed from Table 2, SMDR-IS achieves the highest score in Agg, which demonstrates the suitability of SMDR-IS for advanced underwater vision tasks.
@@ Conclusion
In this study, we propose a novel method for underwater image restoration, namely SMDR-IS, which can proficiently capture multi-scale scene information, by multi-resolution detail extraction with intrinsic supervision, and utilizing the inherent scale-related features in image scenes. To ensure optimal assimilation of information across different scales, we integrated low-resolution inputs by adaptive selective intrinsic supervision within the original-resolution input, thereby amplifying the conveyance of scene information. To mitigate unnecessary interference from irrelevant scenes, we introduced ASISF, which is meticulously designed to regulate the feature propagation process. Furthermore, our multi-degradation loss function strategically guides the network during training. Optimization and constraints at each stage enhance the network’s ability to leverage information at different scales. In the future, we will explore the integration of SMSR-IS into broader computer vision applications, like underwater robotics and autonomous underwater vehicles.
@@ Acknowledgments
This work was supported in part by the Natural Science Foundation of China (Nos. 62301105), National Key Research and Development Program of China (No.2018AAA0100400), China Postdoctoral Science Foundation (NO.2021M701780), and the Cultivation Program for the Excellent Doctoral Dissertation of Dalian Maritime University. We are also sponsored by CAAI-Huawei MindSpore Open Fund and the High Performance Computing Center of Dalian Maritime University.